# Variational inequalities and smooth-fit principle for singular stochastic control problems in Hilbert spaces

### v1

@@ 4. Selecting a specific direction of action: A related optimal stopping problem
Proposition 4.15.
Recall Assumption 2.7, let $x\in H$, $\hat{h}\in K$, and set $h:=\sigma\hat{h}$. If $h\in\mathcal{D}(\mathcal{A})$, then $U$ is differentiable at $x$ along the direction $h$.
Proof.
We assume, without loss of generality, that $U(x)=0$. Indeed, this can be always achieved by a translation. Again, by (i) in Lemma (i) we can write
$U(y)=U_{0}(y)-\frac{\widehat{\kappa}_{1}}{2}|y-x|^{2}_{-1},\quad y\in H,$
for some $\widehat{\kappa}_{1}>0$ and $U_{0}$ convex.
We now argue by contradiction and assume that $U$ (hence $U_{0}$) is not differentiable along the direction $h:=\sigma\hat{h}$, with $\hat{h}\in K$. By convexity of $U_{0}$ (cf. Proposition 4.13) and by Proposition 4.14, there exist $p,\bar{p}\in D^{-}U_{0}(x)\subseteq\mathcal{D}(\mathcal{A})$ such that
$\langle p,h\rangle_{H}<\langle\bar{p},h\rangle_{H}.$
Again by convexity,
$U_{0}(y)\geq\langle p,y-x\rangle_{H}\vee\langle\bar{p},y-x\rangle_{H},\ \ \ \forall y\in H.$
It is then possible to construct $(\widehat{\varphi}_{n})\subseteq\mathcal{X}$ (cf. (3.2)) such that for any $n\in\mathbb{N}$
$\begin{cases}\widehat{\varphi}_{n}(x)=U_{0}(x)=0,\\ \widehat{\varphi}_{n}\leq U_{0},\\ |D\widehat{\varphi}_{n}(x)|_{H}+|\mathcal{A}D\widehat{\varphi}_{n}(x)|_{H}\leq M<\infty,\\ \langle D^{2}\widehat{\varphi}_{n}(x)h,h\rangle_{H}\geq n,\\ \langle D^{2}\widehat{\varphi}_{n}(x)k,\bar{k}\rangle_{H}=0\ \ \ \forall k,\bar{k}\in\mbox{Span}\{h\}^{\perp},\end{cases}$
for some $M>0$. Notice that it follows from the previous properties that
(4.26) $\frac{1}{2}\mbox{Tr}\left[\sigma\sigma^{*}D^{2}\widehat{\varphi}_{n}(x)\right]\geq n.$
As a matter of fact, the sequence
$\widehat{\varphi}_{n}(y)=\frac{1}{2}\langle p+\bar{p},y-x\rangle_{H}+\frac{n}{2}\langle y-x,h\rangle_{H}^{2},\quad n\in\mathbb{N},\,y\in H,$
realizes all the previous requirements, at least in a neighborhood of $x$ (and can be easily modified in order to satisfy them globally on $H$).
Define now $\varphi_{n}(y):=\widehat{\varphi}_{n}(y)-\frac{\widehat{\kappa}_{1}}{2}|y-x|^{2}_{-1}$ for any $y\in H$. Then (cf. (4.26))
$\begin{cases}{\varphi}_{n}(x)=U(x)=0,\\ {\varphi}_{n}\leq U,\\ |D{\varphi}_{n}(x)|_{H}+|\mathcal{A}D{\varphi}_{n}(x)|_{H}\leq M<\infty,\\ \frac{1}{2}\mbox{Tr}\left[\sigma\sigma^{*}D^{2}\varphi_{n}(x)\right]\geq n-\frac{1}{2}\kappa_{1}\mbox{Tr}\left[\sigma\sigma^{*}\mathcal{A}^{-1}\right].\end{cases}$
Because $U$ is a viscosity supersolution at $x$ of (4.19) (cf. Proposition 4.9), one finally has
00 0 $\leq\big{[}\big{(}(\rho-\lambda)-\mathcal{G}\big{)}\varphi_{n}\big{]}(x)=-\frac{1}{2}\mbox{Tr}\left[\sigma\sigma^{*}D^{2}\varphi_{n}(x)\right]-\langle x,\mathcal{A}D\varphi_{n}(x)\rangle_{H}$
$\leq-n+\frac{1}{2}\kappa_{1}\mbox{Tr}\left[\sigma\sigma^{*}\mathcal{A}^{-1}\right]+M|x|_{H}\to-\infty\quad\text{as}\,\,n\to\infty,$
which gives the desired contradiction and thus completes the proof. ∎
@@
The next proposition strenghts the regularity result of Proposition 4.15 in the case
$\overline{\mbox{R}(\sigma)\cap\mathcal{D}(\mathcal{A})}=H,$
where $\mbox{R}(\sigma)$ denotes the range of $\sigma$ (with $\sigma$ satisfying Assumption 2.7).
@@
If
$\overline{\mbox{R}(\sigma)\cap\mathcal{D}(\mathcal{A})}=H,$
then $U$ is (Fréchet) differentiable at $x$, $DU(x)\in\mathcal{D}(\mathcal{A})$, and $DU\in C(H;\mathcal{D}(\mathcal{A}))$, with $\mathcal{D}(\mathcal{A})$ being endowed with the graph norm.
@@
By following the arguments developed at the beginning of Proposition 4.14, we can here assume that $U$ is convex. Then, due to Proposition 4.15 and convexity of $U$, we know that, for each $h\in{R}(\sigma)\cap\mathcal{D}(\mathcal{A})$, the set
$\{\langle h,p\rangle_{H}:\,p\in D^{-}U(x)\}$
is a singleton. Since $\overline{{R}(\sigma)\cap\mathcal{D}(\mathcal{A})}=H$, the set $D^{-}U(x)$ must be a singleton too. Again by convexity of $U$, it follows that $U$ is differentiable at $x\in H$. Furthermore, using Proposition 4.14(i), we conclude that $DU(x)\in\mathcal{D}(\mathcal{A})$.
The last assertion finally follows from the previous conclusions and Proposition 4.14. ∎
@@ 6.1. An irreversible investment problem into energy capacity
where $\Delta$ is the Laplacian operator over $\mathcal{O}$, $\delta>0$ is a depreciation factor, $e(\xi)$ is the initial value of the energy supply at location $\xi$, and $\bf n$ denotes the unitary outer normal vector at the boundary of $\mathcal{O}$; the Neumann boundary condition at $\partial\mathcal{O}$ models the fact that we assume there is no flux of energy at the boundary. Under suitable conditions on the domain, the above PDE is well posed as an abstract evolution equation in $H$. Precisely (see, e.g., [48, Sec.3.1]), defining the operator
@@
The latter controlled SPDE satisfies our Assumptions 2.1, 2.2, and 2.4.
@@ 6.2. An energy balance climate model with human impact
We first describe quickly the basics of an energy balance climate model. The earth’s temperature is taken to be the result of incoming radiation from the sun and outgoing radiation through reflection. We consider temperature on the hemisphere, modeled by the half-circle that we identify with the interval $D:=[-1,1]$, where $\xi\in[-1,1]$ is the sine of latitude. $\mathcal{M}$ is the Borel $\sigma$-algebra on $D$, and $\mu$ is taken to be the Lebesgue measure on $(D,\mathcal{M})$.
@@
and we are back in our setting. In particular, the operator $\mathcal{A}$ with the aforementioned boundary conditions satisfies Assumptions 2.1, 2.2, and 2.4 (for the null Neumann boundary conditions one, see, e.g., [48, Sec.3.1]; for the periodic ones, see Section 5.2 in [28]).
@@
Acknowledgements. Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - Project-ID 317210226 - SFB 1283. Salvatore Federico is grateful to Andrzej Swiech for the valuable discussions on viscosity solutions to variational inequalities in Hilbert spaces that led to the formulation of Remarks 3.9 and 4.10.
Appendix A A result on semiconcave and semiconvex functions
We state here some properties of semiconcave and semiconvex functions on $H$.
Lemma A.1.
  • (i)
    Let $H$ be endowed with an arbitrary norm ||$\cdot$|| ⋅ | and take $F:H\to\mathbb{R}$ be semiconvex (with respect to the norm ||$\cdot$|| ⋅ |); that is, there exists $c_{F}>0$ such that
    $\lambda F(x)+(1-\lambda)F(y)-F(\lambda x+(1-\lambda)y)\geq-\frac{c_{F}}{2}\lambda(1-\lambda)|x-y|^{2},\ \ \forall x,y\in S,\ \lambda\in[0,1].$
    Then
    $x\mapsto F(x)+\frac{c_{F}}{2}|x|^{2}$
    is convex.
  • (ii)
    Let $F:H\to\mathbb{R}$ be both semiconcave and semiconvex (with respect to the norm $|\cdot|_{H}$); that is, there exists $c_{F}>0$ such that
    $\left|\lambda F(x)+(1-\lambda)F(y)-F(\lambda x+(1-\lambda)y)\right|\leq\frac{c_{F}}{2}\lambda(1-\lambda)|x-y|_{H}^{2},$
    for every $x,y\in S$ and $\lambda\in[0,1]$. Then $F\in C^{1,\text{Lip}}(H)$.
Proof.
Proof of (i). This follows from [14, Prop. 1.1.3] (equivalences (a)-(c)), whose proof is not affected by the dimensionality of the space under consideration.
Proof of (ii). This follows by adapting the convex-analytic arguments of the finite-dimensional setting in [14, Ch.3]. We report these here for completeness. Let $x\in S$. Since $F$ is semiconvex and semiconcave, we have that both the supergradient $D^{+}F(x)$ and the subgradient $D^{-}F(x)$ are not empty. We now show that $F$ is differentiable at $x\in H$. Suppose, by the aim of contradiction, that $p_{1}\in D^{-}F(x)$ and $p_{2}\in D^{+}F(x)$ and that $p_{1}\neq p_{2}$. Then, for $h\in H$, by definition of subgradient and by Proposition 3.3.1 in [14, Ch.3] (whose proof does not suffer the dimensionality of the considered space) in the case of linear modulus $\omega(r)=\frac{{c}_{F}}{2}r$, $r\geq 0$, we find
(A-1) $\langle p_{1},h\rangle_{H}-\frac{{c}_{F}}{2}|h|^{2}_{H}\leq F(x+h)-F(x)\leq\langle p_{2},h\rangle_{H}+\frac{{c}_{F}}{2}|h|^{2}_{H}.$
Therefore, we obtain
$0\geq\langle p_{1}-p_{2},h\rangle_{H}-{{c}_{F}}|h|^{2}_{H}.$
Given the arbitrariness of $h$, we can now take $h=\alpha(p_{1}-p_{2})$, for $\alpha>{{c}_{F}}$ so to achieve a contradiction. Hence, $p_{1}=p_{2}$, which, using (A-1), leads to
(A-2) $\frac{\left|F(x+h)-F(x)-\langle p_{1},h\rangle_{H}\right|}{|h|_{H}}\leq\frac{{c}_{F}}{2}|h|_{H},$
and therefore, in particular, to the differentiability of $F$ at $x\in H$ with $DF(x)=p_{1}$.
We now show that $DF:H\to H$ is Lipschitz continuous. To that end, we borrow arguments from the proof of Theorem 3.3.7 in [14, Ch.3], which, once more, does not suffer the dimensionality of the considered space. By (A-2) with $p_{1}=DF(x)$, for any $x,y\in H$ we have
(A-3) $\left|F(y)-F(x)-\langle DF(x),y-x\rangle_{H}\right|\leq\frac{{c}_{F}}{2}|y-x|^{2}_{H}.$
Then, for each $x,v,w\in H$, we have
$\langle DF(x+w),v\rangle_{H}\leq F(x+w+v)-F(x+w)+\frac{{c}_{F}}{2}|v|_{H}^{2},$
and
$\langle DF(x),v\rangle_{H}\geq F(x+v)-F(x)-\frac{{c}_{F}}{2}|v|_{H}^{2}.$
@@ Appendix A A result on semiconcave and semiconvex functions
Hence, from the last two display equations, we obtain
$\langle DF(x+w)-DF(x),v\rangle_{H}\leq F(x+w+v)-F(x+w)-F(x+v)+F(x)+{{c}_{F}}|v|_{H}^{2}$
$=F(x+w+v)-\frac{1}{2}F(x+2v)-\frac{1}{2}F(x+2w)$
$-F(x+w)+\frac{1}{2}F(x+2w)+\frac{1}{2}F(x)$
$-F(x+v)+\frac{1}{2}F(x+2v)+\frac{1}{2}F(x)+{{c}_{F}}|v|_{H}^{2}$
$\leq 2{c}_{F}\left(|w|^{2}_{H}+|v|^{2}_{H}\right),$
where in the last estimate we used the semiconcavity and the semiconvexity of $F$. The latter estimate now implies the claimed Lipschitz property, since
$|DF(x+w)-DF(x)|_{H}=\frac{1}{|w|_{H}}\max_{|v|_{H}=|w|_{H}}\langle DF(x+w)-DF(x),v\rangle_{H}\leq 2{c}_{F}|w|_{H}.$

### v2

@@ 4. Selecting a specific direction of action: A related optimal stopping problem
The last expression tells us that $\hat{\varphi}_{n}\in C^{2}(H)$, and
$D\hat{\varphi}_{n}(y)=\frac{1}{2}\left[\rho_{n}^{\prime}\left(\frac{1}{2}\langle p_{1}-p_{2},y-x\rangle_{H}\right)+1\right](p_{1}+p_{2}).$
Since $p_{1},p_{2}\in\mathcal{D}(\mathcal{A})$, this shows that $\hat{\varphi}_{n}\in\mathcal{X}$. Moreover,
(4.32a) $D\hat{\varphi}_{n}(x)=\frac{1}{2}(p_{1}+p_{2})\in\mathcal{D}(\mathcal{A}),$
(4.32b) $\langle D^{2}\hat{\varphi}_{n}(x)y,z\rangle_{H}=\frac{1}{4}\rho^{\prime\prime}_{n}(0)\langle p_{1}-p_{2},y\rangle_{H}\langle p_{1}-p_{2},z\rangle_{H}\qquad\forall y,z\in H.$
Furthermore, for all $y\in H$,
[…]
Define
$\varphi_{n}(y)\coloneqq\hat{\varphi}_{n}(y)-\frac{\widehat{\kappa}_{1}}{2}|y-x|_{-1}^{2}\qquad\forall y\in H.$
Then $\varphi_{n}\in\mathcal{X}$ and, by (4.32a) and (4.32b), we have
(4.34a) $D\varphi_{n}(x)=\frac{1}{2}(p_{1}+p_{2})-\widehat{\kappa}_{1}\left(\mathcal{A}^{-1}\right)^{2}(y-x).$
Also, by (4.32b) and recalling (4.30f),
(4.35) $\begin{split}\mbox{Tr}\left[\sigma\sigma^{*}D^{2}\varphi_{n}(x)\right]=&\frac{1}{4}n|\sigma^{*}(p_{1}-p_{2})|_{H}^{2}+\widehat{\kappa}_{1}\mbox{Tr}\left[\sigma\sigma^{*}\mathcal{A}^{-1}\right].\end{split}$
By (4.33), $U-\varphi_{n}$ attains a minimum at 0. Because $U$ is a viscosity supersolution at $x$ of (4.19) (cf. Proposition 4.11) and recalling (4.35) and (4.28), one finally has
$0$ $\leq\big[\big((\rho-\lambda)-\mathcal{G}\big)\varphi_{n}\big](x)=-\frac{1}{2}\mbox{Tr}\left[\sigma\sigma^{*}D^{2}\varphi_{n}(x)\right]-\langle x,\mathcal{A}D\varphi_{n}(x)\rangle_{H}$
$\leq-\frac{1}{4}n|\sigma^{*}(p_{1}-p_{2})|_{H}^{2}+\frac{1}{2}\widehat{\kappa}_{1}\mbox{Tr}\left[\sigma\sigma^{*}\mathcal{A}^{-1}\right]+\frac{1}{2}\langle\mathcal{A}(p_{1}+p_{2}),\,x\rangle_{H}\to-\infty\quad\text{as}\,\,n\to\infty,$
which reaches the desired contradiction and thus completes the proof. ∎
@@
Let $\mathrm{R}(\sigma)$ denote the range of $\sigma$. If $\overline{\mathrm{R}(\sigma)}=H,$ then $U$ is (Fréchet) differentiable at all $x\in H$ and $DU(x)\in\mathcal{D}(\mathcal{A})$. Moreover, $DU\in C(H;H)$.
@@
By following the arguments developed at the beginning of Proposition 4.16, we can here assume that $U$ is convex. Then, due to Proposition 4.17 and convexity of $U$, we know that, for each $h\in\mathrm{R}(\sigma)$, the set $\{\langle h,p\rangle_{H}:\,p\in D^{-}U(x)\}$ is a singleton. Since $\overline{\mbox{R}(\sigma)}=H$, this implies that $D^{-}U(x)$ must be a singleton too. Again by convexity of $U$, it follows that $U$ is differentiable at $x\in H$. Furthermore, using Proposition 4.16, we conclude that $DU(x)\in\mathcal{D}(\mathcal{A})$.
Finally, the last assertion follows from the Corollary at p. 20 in [59]. ∎
@@ 6.1. An irreversible investment problem into energy capacity
where $\Delta$ is the Laplacian operator over $\mathcal{O}$, $\delta>0$ is a depreciation factor, $e(\xi)$ is the initial value of the energy supply at location $\xi$, and $\bf n$ denotes the unitary outer normal vector at the boundary of $\mathcal{O}$; the Neumann boundary condition at $\partial\mathcal{O}$ models the fact that we assume there is no flux of energy at the boundary. Under suitable conditions on the domain, the above PDE is well posed as an abstract evolution equation in $H$. Precisely (see, e.g., [50, Sec.3.1]), defining the operator
@@
The latter controlled SPDE satisfies our Assumptions 2.1 and $2.$
@@ 6.2. An energy balance climate model with human impact
We first describe quickly the basics of an energy balance climate model. The earth’s temperature is taken to be the result of incoming radiation from the sun and outgoing radiation through reflection. We consider temperature on the hemisphere, modeled by the half-circle that we identify with the interval D:=[1,1]D:=[-1,1], where $\xi\in[-1,1]$ is the sine of latitude. $\mathcal{M}$ is the Borel $\sigma$-algebra on $D$, and $\mu$ is taken to be the Lebesgue measure on $(D,\mathcal{M})$.
@@
and we are back in our setting. In particular, the operator $\mathcal{A}$ with the aforementioned boundary conditions satisfies Assumptions 2.1 and 2.2 (for the null Neumann boundary conditions one, see, e.g., [50, Sec.3.1]; for the periodic ones, see Section 5.2 in [29]).
@@
Acknowledgements. Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - Project-ID 317210226 - SFB 1283. Salvatore Federico is grateful to Andrzej Swiech for the valuable discussions on viscosity solutions to variational inequalities in Hilbert spaces that led to the formulation of Remarks 3.10 and 4.12.
Appendix A Proof of Proposition 3.5
We limit ourselves to provide a sketch of the proof, given that this follows closely from arguments used in [28, Ch. 2 and Sec. 3.6.2] (for the infinite-dimensional case with regular controls) and in [22] (for the finite-dimensional case with singular controls).
To this end, we let $t\geq 0$, we take a cylindrical Wiener process $(W_{s})_{s\in[t,\infty)}$ on $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ such that $W_{t}=0$ a.s., and, for $s\geq t$, we set $\mathcal{F}^{t,0}_{s}:=\sigma(W_{r},\,t\leq r\leq s)$ and denote by $\mathcal{F}^{t}_{s}$ its completion with $\mathbb{P}$-null sets. We then write $\mathbb{F}^{t,0}:=(\mathcal{F}^{t,0}_{s})_{s\in[t,\infty)}$ and $\mathbb{F}^{t}$ for the augmented filtration.
In analogy with (2.18) and (2.19), recalling (2.17), we let
(A-1) $\mathcal{S}_{t}$ $:=$ $\{\nu:\Omega\times[t,\infty)\to[0,\infty):\,\text{$\nu_{\cdot}$ is}\,\,\mathbb{F}^{t}-\text{adapted and such that}\,\,s\mapsto\nu_{s}$
[…]
In the sequel, we set $\nu_{t^{-}}:=0$ for any $\nu\in\mathcal{S}_{t}$ and define
(A-2) $\mathcal{I}_{t}$ $:=$ $\Big\{(\vartheta,\nu):\Omega\times[t,\infty)\to\Theta\times[0,\infty):\ \vartheta_{\cdot}\ \mbox{is}\ \mathbb{F}^{t}-\mbox{adapted},\,\,\nu_{\cdot}\in\mathcal{S}_{t}\,\,\,\text{and}$
$\mathsf{E}\bigg[\Big|\int_{t^{-}}^{T}|{\vartheta}_{s}|_{H}\,\mathrm{d}\nu_{s}\Big|^{2}\bigg]+\mathsf{E}\bigg[\int_{t^{-}}^{T}|{\vartheta}_{s}|^{2}_{H}\,\mathrm{d}\nu_{s}\bigg]<\infty\,\,\forall T>0\Big\}.$
Finally, we set (cf. (A-3))
(A-3) $\hat{\mathcal{I}}_{t}:=\left\{I=(\vartheta,\nu)\in\mathcal{I}_{t}:\ \mathsf{E}\bigg[\int_{t^{-}}^{\infty}e^{-\rho(s-t)}\mathrm{d}\nu_{s}\bigg]<\infty\right\}.$
Now, for $I:=(\vartheta,\nu)\in\hat{\mathcal{I}_{t}}$ and $x\in H$, we let $(X^{t,x,I})_{s\in[t,\infty)}$ be the unique mild solution to
(A-4) $\mathrm{d}X^{I}_{s}=\mathcal{A}X^{I}_{s}\mathrm{d}t+\sigma\mathrm{d}W_{s}+\mathrm{d}I_{s},\quad s\geq t,\ \ \ X_{t^{-}}^{I}=x\in H;$
that is,
(A-5) $X_{s}^{t,x,I}=e^{(s-t)\mathcal{A}}x+\int_{t^{-}}^{s}e^{(s-r)\mathcal{A}}\sigma\mathrm{d}W_{r}+\int_{t^{-}}^{s}e^{(s-r)\mathcal{A}}\mathrm{d}I_{r},\quad s\geq t.$
For $u\in[0,s]$, it is then easy to observe from direct manipulations on (A-5), that the following flow property holds:
(A-6) $X_{s}^{0,x,I}=e^{(s-u)\mathcal{A}}X_{u}^{0,x,I}+\int_{u^{-}}^{s}e^{(s-r)\mathcal{A}}\sigma\mathrm{d}W_{r}+\int_{u^{-}}^{s}e^{(s-r)\mathcal{A}}\mathrm{d}I_{r}=X_{s}^{u,X_{u}^{t,x,I},I}.$
Step 1. We first provide a sketch of the proof of a weaker version of (3.19), namely for any time $\eta\in[0,\infty)$ it holds
(A-7) $V(x)=\inf_{I\in\hat{\mathcal{I}_{0}}}\mathsf{E}\left[\int_{0^{-}}^{\eta}e^{-\rho t}\left(G(X^{x,I}_{t})\mathrm{d}t+\mathrm{d}\nu_{t}\right)+e^{-\rho\eta}V(X_{\eta}^{x,I})\right],\quad x\in H.$
Proof of $\geq$” in (A-7). As in [22, Lemma 5.1] (see [28, Lemma 1.99] for the infinite-dimensional setting with regular controls), now denote by $\overline{I}$ the $\mathbb{F}^{t,0}$-predictable process such that $\overline{I}\in\hat{\mathcal{I}_{t}}$ and $\overline{I}=I$ $\mathbb{P}\otimes\mathrm{d}t$-a.e. on $\Omega\times[t,\infty)$. Accordingly, we set $\overline{X}_{s}^{t,x,\overline{I}}$ to be given by (A-5) with $\overline{I}$ instead of $I$. Then, using that $\overline{X}_{s}^{t,x,\overline{I}}$ and ${X}_{s}^{t,x,{I}}$ are indistinguishable and exploiting the flow property (A-6) (now for the mild solution $\overline{X}$), we obtain
(A-8) $\hskip 56.9055pt\mathsf{E}\bigg[\int_{0^{-}}^{\infty}e^{-\rho t}\Big({G}(X_{t}^{0,x,I})\mathrm{d}t+\mathrm{d}\nu_{t}\Big)\bigg]$
$=\mathsf{E}\bigg[\int_{0^{-}}^{\eta}e^{-\rho t}\Big({G}({X}_{t}^{0,x,I})\mathrm{d}t+\mathrm{d}{\nu}_{t}\Big)\bigg]+\mathsf{E}\bigg[\int_{\eta^{-}}^{\infty}e^{-\rho t}\Big({G}(\overline{X}_{t}^{0,x,\overline{I}})\mathrm{d}t+\mathrm{d}\overline{\nu}_{t}\Big)\bigg]$
$=\mathsf{E}\bigg[\int_{0^{-}}^{\eta}e^{-\rho t}\Big({G}({X}_{t}^{0,x,I})\mathrm{d}t+\mathrm{d}{\nu}_{t}\Big)\bigg]+\mathsf{E}\bigg[e^{-\rho\eta}\mathsf{E}\bigg[\int_{\eta^{-}}^{\infty}e^{-\rho(t-\eta)}\Big({G}(\overline{X}_{t}^{0,x,\overline{I}})\mathrm{d}t+\mathrm{d}\overline{\nu}_{t}\Big)\Big|\,\mathcal{F}^{0,0}_{\eta}\bigg]\bigg]$
$=\mathsf{E}\bigg[\int_{0^{-}}^{\eta}e^{-\rho t}\Big({G}({X}_{t}^{0,x,I})\mathrm{d}t+\mathrm{d}{\nu}_{t}\Big)\bigg]+\mathsf{E}\bigg[e^{-\rho\eta}\mathsf{E}\bigg[\int_{\eta^{-}}^{\infty}e^{-\rho(t-\eta)}\Big({G}(X_{t}^{\eta,X_{\eta}^{0,x,\overline{I}},\overline{I}})\mathrm{d}t+\mathrm{d}\overline{\nu}_{t}\Big)\Big|\,\mathcal{F}^{0,0}_{\eta}\bigg]\bigg].$
In order to take care of the conditional expectation in the last display equation, we proceed as in Equation (2.29) in [28] (see also Step 1 in the proof of [22, Thm. 4.2]). In particular, we define the regular conditional probability on $(\Omega,\mathcal{F})$ given by $\mathbb{P}_{\omega}(A):=\mathbb{P}(A\,|\,\mathcal{F}^{0,0}_{\eta})$ and for any $A\in\mathcal{F}$ $\mathbb{P}$-a.e. $\omega$, and we denote by $\mathbb{E}_{\omega}$ its expectation. Moreover, we introduce the control system
$I^{\omega}(\cdot):=(\Omega,\mathcal{F}_{\omega},\mathbb{P}_{\omega},\mathbb{F}^{\omega,\overline{\tau}_{\varepsilon}}:=(\mathcal{F}^{\omega,\eta}_{s})_{s\in[\eta,\infty)},(W_{s}-W_{\eta})_{s\in[\eta,\infty)},\overline{I}|_{[\eta^{-},\infty)}(\cdot))\in\hat{\mathcal{I}}_{\eta},$
where $\mathcal{F}^{\omega}$ is the completion of $\mathcal{F}$ with $\mathbb{P}^{\omega}$-null sets and $\mathbb{F}^{\omega,\eta}$ the augmentation with $\mathbb{P}^{\omega}$-null sets of the sigma-algebra generated by the increments $W_{\cdot}-W_{\eta}$,
Then, appealing to standard properties of regular conditional probabilities (see the remarks at Page 102 in [28]), law-invariance property of the state process (see also [28, Hypothesis 2.29]), and recalling the definition of the cost functional (2.21), we can write
$\mathsf{E}\bigg[\int_{\eta^{-}}^{\infty}e^{-\rho(t-\eta)}\Big({G}(X_{t}^{\eta,X_{\eta}^{0,x,\overline{I}},\overline{I}})\mathrm{d}t+\mathrm{d}\overline{\nu}_{t}\Big)\Big|\,\mathcal{F}^{0,0}_{\eta}\bigg]$
$=\mathsf{E}_{\omega}\bigg[\int_{\eta^{-}}^{\infty}e^{-\rho(t-\eta)}\Big({G}(X_{t}^{\eta,X_{\eta}^{0,x,\overline{I}},\overline{I}})\mathrm{d}t+\mathrm{d}\overline{\nu}_{t}\Big)\bigg]=\mathcal{J}(\overline{X}_{\eta}^{0,x,{I}}(\omega);I^{\omega})\geq V(\overline{X}_{\eta}^{0,x,{I}}(\omega)).$
@@
Hence, using this last display equation into (A-8) we obtain
$\mathsf{E}\bigg[\int_{0^{-}}^{\infty}e^{-\rho t}\Big({G}(X_{t}^{0,x,I})\mathrm{d}t+\mathrm{d}\nu_{t}\Big)\bigg]$
(A-9) $\geq\mathsf{E}\bigg[\int_{0^{-}}^{\eta}e^{-\rho t}\Big({G}({X}_{t}^{0,x,I})\mathrm{d}t+\mathrm{d}{\nu}_{t}\Big)+e^{-\rho\eta}V(\overline{X}_{\eta}^{0,x,{I}})\bigg]$
By taking the infimum over all $I\in\hat{\mathcal{I}_{0}}$ it thus follows that
(A-10) $V(x)\geq\inf_{I:=(\vartheta,\nu)\in\hat{\mathcal{I}_{0}}}\mathsf{E}\left[\int_{0^{-}}^{\eta}e^{-\rho t}\left(G(X^{x,I}_{t})\mathrm{d}t+\mathrm{d}\nu_{t}\right)+e^{-\rho\eta}V(X_{\eta}^{x,I})\right].$
Proof of $\leq$ in (A-7). Given the continuity of $V$ (see Proposition 3.1) and the continuity of the control functional $\mathcal{J}(\cdot\,;I)$ (due to semiconcavity, see the proof of item (iv) in the proof of Proposition 3.1), one can argue as in Part 2 of the proof of [28, Thm. 2.24] (see also Step 2 in the proof of [22, Thm. 4.2] for the finite-dimensional setting with singular controls) to conclude.
Step 2. Having the weaker version (A-7) and given the continuity of $V$ implied by its semiconcavity, one can then argue as in the proof of [28, Th. 3.70] (whose arguments still hold in the case of singular controls) to conclude the validity of (3.19).